A 1.5 micro C charge is 14 mm north of location X. A -2.3 Micro C charge is 11 mm south of X.

A) What is the electric potential at location X? Assume the potential is zero at infinity.

B) How much electric potential energy would be added to the system if a 1.7 micro C charge were placed at location X.

Show all your Work.
Answer #1

let Q1 = 1.5 micro C = 1.5*10^-6 C
r1 = 14 mm = 0.014 mm

Q2 = -2.3 micro C = -2.3*10^-6 C
r2 = 11 mm = 0.011 m

A) potential at location X, V = V1 + V2

= k*Q1/r1 + k*Q2/r2

= 9*10^9*1.5*10^-6/0.014 + 9*10^9*(-2.3*10^-6)/0.011

= -9.18*10^5 V

B) electric potential energy added = change in potential energy of the charge system

= Uf - Ui

= q*(Vf - Vi)

= 1.7*10^-6*(-9.18*10^5 - 0 )

= -1.56 J